find the present value and the compound discount of 6600$ due in 7 years 6 months. if interest is 7.2% quarterly.

Answer by addingup(3677)

I think the problem should say interest of 7.2% per annum compounded quarterly. But that's not what you say, you say 7.2% quarterly, so here it goes:
.
PV = (Initial cash flow)/(1+rate)^n
7 years 6 months = 30 quarters
PV = 6600/(1+0.072)^30 = 820
-.-.-.-.-.-.-.-
Note: if it was 7.2%/year compounded quarterly, it would be:
.
7 years 6 months = 30 quarters
Interest: 0.072/4 = 0.018
PV = 6600/(1+0.018)^30 = 3866.45
.
I don't understand what you mean by "the compound discount". You compound when you go from Present Value to Future Value and discount when you go from Future to Present